OpenStudy (anonymous):

The velocity of the function (in meters per second) is given for a particle moving along a line. Find (a) the displacement and (b) the distance traveled by the particle during the given time interval. v(t)=t^2-2t-8, [1,6]

OpenStudy (nincompoop):

understand that velocity is a derivative of a position (displacement) function

OpenStudy (anonymous):

So first I took \[\int\limits_{1}^{6}t^2-2t-8dt\]

OpenStudy (anonymous):

why?

OpenStudy (nincompoop):

v(t) = lim ∆t -> 0= [x(t+Δt) - t]÷∆t = dx/dt

OpenStudy (anonymous):

I don't know that form. We're supposed to use the Fundamental Theorem of Calculus.

OpenStudy (nincompoop):

alright, let us assume that your integral set up is correct integrate your function then apply V(b) - V(a)

OpenStudy (anonymous):

for v(b)-v(a)=v(6)-v(1)=-10/3

OpenStudy (nincompoop):

oh it's minus

OpenStudy (nincompoop):

\[\frac{ t^3 }{ 3 }- \frac{ 2t^2 }{ 2 } - 8t ] _{1}^{6}\]

OpenStudy (anonymous):

I actually think I figured out the problem.

OpenStudy (nincompoop):

[6^2/3 - 6^2 - 8(6)] - [(1/3-1-8)]

OpenStudy (anonymous):

I figured it out. Thank you for helping though! I just needed to take some time to try and work it out.